OpenStudy (anonymous):

solve the system of equations below 4x+y=16 and 2x+3y=-2 A) (4,-5) B) (-5,4) C) (5,4) D) (5,-4)

OpenStudy (anonymous):

please help! anyone! haha

OpenStudy (anonymous):

@Damiselchild can you help me? :)

OpenStudy (anonymous):

@Littlebird could you help me out?

jimthompson5910 (jim_thompson5910):

are you able to isolate y in 4x+y=16 ?

OpenStudy (anonymous):

would it be y=4x+16?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

you subtract 4x from both sides though

jimthompson5910 (jim_thompson5910):

to get y = -4x + 16

jimthompson5910 (jim_thompson5910):

move onto the second equation 2x+3y=-2 2x+3(-4x+16)=-2 ... replace y with -4x+16 solve for x

OpenStudy (anonymous):

@jim_thompson5910 x=5

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

y = -4x + 16 y = -4(5) + 16 ... replace x with 5 (since we now know that x = 5) y = ??

OpenStudy (anonymous):

y=-4 so its (5,-4)

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

if you were to graph the original equations, you'd see that they cross at (5,-4)

jimthompson5910 (jim_thompson5910):

and it's a good idea to check the solution in both 4x+y=16 and 2x+3y=-2

OpenStudy (littlebird):

Here is a site that can summarize jim_thompson's teachings www.webmath.com/solver2.html